You want to find three consecutive integers whose sum is 84.
A. Why does the equation n+(n+1)+(n+2)=84 model the situation ?
B. solve the equation in part (a) then find the three consecutive integers

Found 2 solutions by oscargut, stanbon:
Answer by oscargut(2103)

a) Yes the equation is right since n,n+1 and n+2 are consecutive integers
(n is the smallest integer)
b) n+(n+1)+(n+2)=84
3n+3 = 84
3n = 81
n = 27
Integers are: 27,28 and 29


If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks

Answer by stanbon(75887)

You want to find three consecutive integers whose sum is 84.
A. Why does the equation n+(n+1)+(n+2)=84 model the situation ?
1st integer: n
2nd integer: n+1
3rd integer: n+1+1 = n+2
===================================
B. solve the equation in part (a) then find the three consecutive integers
3n+3 = 84
3n = 81
n = 27
n+1 = 28
n+2 = 29
===============
Cheers,
Stan H.
================